OpenStudy (anonymous):

If tan^4 theta+tan^2 theta=1 then cos^4theta+cos^2 theta is

OpenStudy (anonymous):

tan^2 theta(tan^2 theta+1) sin^2/cos^2(sin^2/cos^2+1)

OpenStudy (anonymous):

sin^2 theta/cos^2 theta(1/cos^2 theta)

OpenStudy (anonymous):

(1-cos^2 theta)/cos^2 theta*1/(cos^2theta)

OpenStudy (anonymous):

\[(1/coa^2 \theta-\cos^2 \theta/\cos^2 \theta)*1/\cos^2 \theta\]

OpenStudy (anonymous):

\[(1-\cos^2 \theta)/\cos^4 \theta=0\]

OpenStudy (dumbcow):

** should equal 1

OpenStudy (anonymous):

so \[\cos^4 \theta+\cos^2\theta=1\]

OpenStudy (anonymous):

my steps are correct ahh.i got 1 as the answer is it right?

OpenStudy (dumbcow):

yes ... seems like you dont need any help :)

OpenStudy (anonymous):

Thank you